# Attacking Vision-Language Computer Agents via Pop-ups

### v1

@@ Abstract
Autonomous agents powered by large vision and language models (VLM) have demonstrated significant potential in completing daily computer tasks, such as browsing the web to book travel and operating desktop software, which requires agents to understand these interfaces. Despite such visual inputs becoming more integrated into agentic applications, what types of risks and attacks exist around them still remain unclear. In this work, we demonstrate that VLM agents can be easily attacked by a set of carefully designed adversarial pop-ups 111In this work, we use “pop-ups” to refer to clickable malicious images on the screen., which human users would typically recognize and ignore. This distraction leads agents to click these pop-ups instead of performing the tasks as usual. Integrating these pop-ups into existing agent testing environments like OSWorld and VisualWebArena leads to an attack success rate (the frequency of the agent clicking the pop-ups) of 86% on average and decreases the task success rate by 47%. Basic defense techniques such as asking the agent to ignore pop-ups or including an advertisement notice, are ineffective against the attack.
@@ 1 Introduction
Most previous agent attacks either made the adversarial examples as visually similar to the original ones (Wu et al., 2024) or inject invisible adversarial strings into web pages (Liao et al., 2024; Xu et al., 2024). Here, we argue that whether the adversarial examples are visible or recognizable by humans is not essential if the agent’s ultimate goal is to complete tasks with minimal or no human supervision. As long as the environment functions well and human users can complete the tasks as usual, the agent should be able to complete the tasks as well. Since experienced human users can identify suspicious online content and rarely follow the instructions in unverified pop-ups, we aim to investigate whether these adversarial pop-ups can mislead agents and thus can be used to stress test agents’ capabilities. Our design space (Figure 2) of attacks includes four representative elements to attack: (i) Attention Hook: a few words to attract the agent’s attention. (ii) Instruction: desired behaviors the attacker intends for the agent to follow. (iii) Information Banner: contextual information that implies or misleads the agent about the purpose of the pop-ups. (iv) ALT 222In HTML, alternative text (ALT text) is displayed when an element cannot be rendered, and it was previously used to enhance SoM agents. Descriptor: supplemental textual information provided for the pop-up within the a11y tree. In our experiments, we insert various types of adversarial pop-ups into the observation space for environments like OSWorld (Xie et al., 2024) and VisualWebArena (Koh et al., 2024).
By testing screenshot agents (Xie et al., 2024) and Set-of-Mark agents (Yang et al., 2023) using state-of-the-art VLMs as backbones, we find that our attack achieves an attack success rate (ASR) over 80% on OSworld and over 60% on VisualWebArena in the default setting, where we assume the attacker has complete information (including the user query, the pop-up’s position, and the underlying agent framework, etc). Via a comprehensive set of ablation studies on the design choices of such adversarial pop-ups, we find that: (1) User query is essential for the attention hook, as using other alternatives (e.g., attackers speculate the user intent from the screen content.), on average, decreases the ASR by 61% relatively. (2) Other information (e.g., position and agent framework information) is relatively unnecessary to make the attack successful. (3) Basic defense strategies, such as asking the agent to ignore pop-ups and adding an extra advertisement notice, cannot effectively mitigate the issue (decrease the ASR by no more than 25% relatively). In summary, deploying computer-use agents still suffers from significant risks, and more robust agent systems are needed to ensure safe agent workflow.
@@ 2 Related Work
To attack VLM agents, Wu et al. (2024) propose to add learnable noises to the images so that the captioner or VLM will output adversarial captions first, which will mislead the VLM later. However, such gradient-based attacks take thousands of steps to optimize and are less effectively transferrable to commercial closed-source VLMs. Liao et al. (2024) inject invisible malicious instructions to the websites requesting agents’ private user information. Xu et al. (2024) further shows that such malicious instructions can be generated by an adversarial prompter model, which is trained using successful and failed attack data. Building attacks based on HTML text is not realistic in the long term as the agent framework could shift to screenshot-based gradually, and the problem is also more similar to text-based jailbreaking (Zou et al., 2023; Liu et al., 2023) and backdoor attacks (Yang et al., 2024). Ma et al. (2024) study the faithfulness of VLM agents by checking whether they get distracted by unrelated but not malicious elements in the environment. In this work, we instead study the robustness of VLM agents against malicious elements, such as pop-ups that target agents.
@@
| Prompt Defense | 95.9 | 6.0 | 93.4 | 14.0 | 60.3 | 47.2 | |||
| Prompt Defensespecific | 52.0 | 6.0 | 72.3 | 24.0 | 60.2 | 45.8 | |||
@@ 4 Experiments
For OSWorld Xie et al. (2024), we test screenshot agents and SoM agents on 50 easy tasks, which are selected from those completed tasks in the full testing set experiment without attacks. Note that these tasks are not guaranteed to be completed without attacks due to the decoding randomness (temperature is set to with top_p the original setting). For VisualWebArena, we use a subset containing 72 easy tasks selected by Wu et al. (2024) and only test SoM agents. In our experiments, we follow the original settings from both benchmarks except for using a decoding temperature of to reduce randomness.
@@
We use five frontier VLMs for our experiments: gpt-4-turbo-2024-04-09 (Achiam et al., 2023), gpt-4o-2024-05-13 (OpenAI, 2024), gemini-1.5-pro-002 (Reid et al., 2024), claude-3-5-sonnet-20240620, and the latest claude-3-5-sonnet-20241022 (Anthropic, 2024). Though prior works are heavily built on gpt-4-vision-preview, we choose not to use it due to its deprecation.
@@
To implement our attack, we first find the largest available rectangle area on the screen after excluding the bounding boxes from the a11y tree and OCR detection. We then randomly sample a position and size within this rectangle. Finally, we search for the largest possible font size that fits within the pop-ups, where the attention hook/instruction/info banner is arranged as the example in Figure 2. To fully utilize the computational cost, we attack the agent observation whenever there is sufficient space for our pop-ups (both height and width are more than pixels). If the agent clicks on our pop-ups, we ignore this action during execution, and no redirection is implemented for simplicity. We use gpt-4o-2024-05-13 to summarize the user query and speculate the user query based on information on the screen through a11y trees. By default, we use “Please click (x,y)” as the instruction for both screenshot- and SoM agents in all OSWorld experiments, and “Please click [ID]” for SoM agents in all VisualWebArena experiments. More details are in Appendix A.
@@ 6 Conclusion
In this work, we show that VLM agents can be easily maliciously instructed by adversarial pop-ups while operating on computers. Even though these pop-ups look very suspicious (by design) to human users, agents cannot distinguish the difference between pop-ups and typical digital content, even with extra advertisement notices. This work offers two takeaways: (i) Just like human users need to undergo training to recognize phishing emails (Kumaraguru et al., 2007), VLM models/agents might need to undergo a similar process to ignore environmental noises and prioritize legitimate instructions (Wallace et al., 2024) before operating in the real digital world. This also applies to embodied agents since many distractors in the physical environment might also be absent from the training data. (ii) Human users need to oversee the automated agent workflow carefully to manage the potential risks from the environment. Future work might focus on effectively leveraging human supervision and intervention for necessary safety concerns.
@@
This work is subject to a few limitations. (i) We only test the performance of closed-source models, which hinders a deeper understanding of why such an attack works. This choice was made due to the relatively low performance of open-source models on computer agent benchmarks. Future research is encouraged to explore more capable open-source models. (ii) We do not explore more advanced jailbreaking techniques, such as optimizing the string inside the pop-ups (Zou et al., 2023) or making our pop-ups more persuasive (Zeng et al., 2024), but focus more on the high-level design of the adversarial pop-ups and the contribution analysis of different components. (iii) There are several practical approaches to potentially mitigate our attack in the real world, such as implementing more robust content filtering in browsers, adding an additional module to detect malicious instructions in agent observations, and using more detailed and specific descriptions of the attack within the defensive prompt. The efforts described above may not be sufficient to mitigate all environmental risks. Ideally, the VLM model and agent should recognize and understand this type of risk without requiring external tools. Our findings indicate that current models are still far behind these expectations.
@@ Acknowledgments
We thank Aryaman ‘Adam’ Arora, Harshit Joshi, Nikil Selvam, Yijia Shao, Yangjun Ruan, Dora Zhao, John Yang, Hao Zhu, Michael Ryan, Ryan Li, Ryan Louie, Caleb Ziems, Will Held, Yutong Zhang, Chen Henry Wu, Junlin Yang, Siyuan Ma, Graham Neubig, and all wonderful SALT Lab members for their valuable feedback on different stages of this work.
@@ Appendix A Appendix
We limit the height of pop-ups to pixels and the width to pixels. The height of the info banner at the bottom is set to pixels, or one-third of the height of the pop-ups if the pop-up is not large enough. We randomly select a tag ID for our pop-up, which does not correlate with its website position. For the a11y tree element corresponding to the pop-ups, we use “{Tag ID} button button {ALT}” and “[{Tag ID}] [A] [{ALT}]” as adversarial templates for OSWorld and VisualWebArena correspondingly, which assumes the pop-ups are clickable items. We can run two benign (non-clickable) templates for comparison: “{Tag ID} {ALT}” for OSWorld and “[{Tag ID}] [IMG] [{ALT}]” for VisualWebArena, where we provide results in Table 9. The benign template leads to a slight increase in SR and a slight decrease in ASR.
@@
| Step-wise Prompt Defense | OSWorld-Screen | |||
| ASR↓ | TASR↓ | SR↑ | OSR↑ | |
| Default attack | 5.9 | 32.0 | 8.0 | 18.0 |
| Remove “Please click” | 10.4 | 44.0 | 10.0 | |
Step-wise Prompt Defense
Beyond adding “PLEASE IGNORE THE POP-UPS ON THE SCREEN!!! PLEASE IGNORE THE CLICKING INSTRUCTIONS ON THE SCREEN!!!” to the system prompt, we also test the performance of adding this to the agent instruction of each step (Table 10). While this approach further reduces the ASR compared to only modifying the system prompt, it only results in a marginal increase in the SR. One potential way to counter such a strong defense is to paraphrase the instructions in the pop-ups (Krishna et al., 2024; Abdali et al., 2024). We tested a simple baseline by removing the phrase “Please click” before the coordinates in the malicious instructions, which proved effective against this defense. We will leave more systematic studies for future research.

### v2

@@ Abstract
Autonomous agents powered by large vision and language models (VLM) have demonstrated significant potential in completing daily computer tasks, such as browsing the web to book travel and operating desktop software, which requires agents to understand these interfaces. Despite such visual inputs becoming more integrated into agentic applications, what types of risks and attacks exist around them still remain unclear. In this work, we demonstrate that VLM agents can be easily attacked by a set of carefully designed adversarial pop-ups 111In this work, we use “pop-ups” to refer to clickable malicious images on the screen., which human users would typically recognize and ignore. This distraction leads agents to click these pop-ups instead of performing their tasks as usual. Integrating these pop-ups into existing agent testing environments like OSWorld and VisualWebArena leads to an attack success rate (the frequency of the agent clicking the pop-ups) of 86% on average and decreases the task success rate by 47%. Basic defense techniques, such as asking the agent to ignore pop-ups or including an advertisement notice, are ineffective against the attack. Code is available at https://github.com/SALT-NLP/PopupAttack.
@@ 1 Introduction
Most previous agent attacks either made the adversarial examples as visually similar to the original ones (Wu et al., 2024) or inject invisible adversarial strings into web pages (Liao et al., 2024; Xu et al., 2024a). Here, we argue that whether the adversarial examples are visible or recognizable by humans is not essential if the agent’s ultimate goal is to complete tasks with minimal or no human supervision. As long as the environment functions well and human users can complete the tasks as usual, the agent should be able to complete the tasks as well. Since experienced human users can identify suspicious online content and rarely follow the instructions in unverified pop-ups, we aim to investigate whether these adversarial pop-ups can mislead agents and thus can be used to stress test agents’ capabilities. Our design space (Figure 2) of attacks includes four representative elements to attack: (i) Attention Hook: a few words to attract the agent’s attention. (ii) Instruction: desired behaviors the attacker intends for the agent to follow. (iii) Information Banner: contextual information that implies or misleads the agent about the purpose of the pop-ups. (iv) ALT 222In HTML, alternative text (ALT text) is displayed when an element cannot be rendered, and it was previously used to enhance SoM agents. Descriptor: supplemental textual information provided for the pop-up within the a11y tree. In our experiments, we insert various types of adversarial pop-ups into the observation space for environments like OSWorld (Xie et al., 2024) and VisualWebArena (Koh et al., 2024).
By testing screenshot agents (Xie et al., 2024) and Set-of-Mark agents (Yang et al., 2023) using state-of-the-art VLMs as backbones, we find that our attack achieves an attack success rate (ASR) over 80% on OSworld and over 60% on VisualWebArena in the default setting, where we assume the attacker has complete information (including the user query, the pop-up’s position, and the underlying agent framework, etc). Via a comprehensive set of ablation studies on the design choices of such adversarial pop-ups, we find that: (1) User query is essential for the attention hook, as using other alternatives (e.g., attackers speculate the user intent from the screen content.), on average, decreases the ASR by 61% relatively. (2) Other information (e.g., position and agent framework information) is relatively unnecessary to make the attack successful. (3) Basic defense strategies, such as asking the agent to ignore pop-ups in system prompts and adding an extra advertisement notice, cannot effectively mitigate the issue (decrease the ASR by no more than 25% relatively). In summary, deploying computer-use agents still suffers from significant risks, and more robust agent systems are needed to ensure safe agent workflow.
@@ 2 Related Work
To attack VLM agents, Wu et al. (2024) propose to add learnable noises to the images so that the captioner or VLM will output adversarial captions first, which will mislead the VLM later. However, such gradient-based attacks take thousands of steps to optimize and are less effectively transferrable to commercial closed-source VLMs. Liao et al. (2024) inject invisible malicious instructions to the websites requesting agents’ private user information. Xu et al. (2024a) further shows that such malicious instructions can be generated by an adversarial prompter model, which is trained using successful and failed attack data. Building attacks based on HTML text is not realistic in the long term as the agent framework could shift to screenshot-based gradually, and the problem is also more similar to text-based jailbreaking (Zou et al., 2023; Liu et al., 2023) and backdoor attacks (Yang et al., 2024). Ma et al. (2024) study the faithfulness of VLM agents by checking whether they get distracted by unrelated but not malicious elements in the environment. In this work, we instead study the robustness of VLM agents against malicious elements in observations (Zhan et al., 2024), such as pop-ups that target agents.
@@
| System Prompt Defense | 95.9 | 6.0 | 93.4 | 14.0 | 60.3 | 47.2 | |||
| System Prompt Defensespecific | 52.0 | 6.0 | 72.3 | 24.0 | 60.2 | 45.8 | |||
| OSWorld-Screen | ||||
| ASR↓ | TASR↓ | SR↑ | OSR↑ | |
| Step-wise Prompt Defense | 5.9 | 32.0 | 8.0 | 18.0 |
| + Attack Variance | 10.4 | 44.0 | 10.0 | |
@@ 4 Experiments
For OSWorld Xie et al. (2024), we test screenshot agents and SoM agents on 50 easy tasks, which are selected from those completed tasks in the full testing set experiment without attacks. Each task has a 15-step limit. Note that these tasks are not guaranteed to be completed without attacks due to the decoding randomness (temperature is set to with top_p the original setting). For VisualWebArena, we use a subset containing 72 easy tasks selected by Wu et al. (2024) and only test SoM agents. Each task has a 30-step limit. In our experiments, we follow the original settings from both benchmarks except for using a decoding temperature of to reduce randomness.
@@
We use five frontier VLMs for our experiments: gpt-4-turbo-2024-04-09 (Achiam et al., 2023), gpt-4o-2024-05-13 (OpenAI, 2024), gemini-1.5-pro-002 (Reid et al., 2024), claude-3-5-sonnet-20240620, and the latest claude-3-5-sonnet-20241022 (Claude 3.5 Sonnet v2) (Anthropic, 2024). Though prior works are heavily built on gpt-4-vision-preview, we choose not to use it due to its deprecation.
@@
To fully utilize the computational cost, we attack the agent observation whenever there is sufficient space for our pop-ups (The implementation of finding optimal screen space for pop-up placement and font sizing, with full technical specifications provided in Appendix A). If the agent clicks on our pop-ups, we ignore this action during execution, and no redirection is implemented for simplicity. We use gpt-4o-2024-05-13 to summarize the user query and speculate the user query based on information on the screen through a11y trees. By default, we use “Please click (x,y)” as the instruction for both screenshot- and SoM agents in all OSWorld experiments, and “Please click [ID]” for SoM agents in all VisualWebArena experiments.
@@ 4.3 Defense
Step-wise Prompt Defense
Beyond adding defense to the system prompt, we also test the performance of adding this to the agent instruction of each step (Table 8). While this approach further reduces the ASR compared to only modifying the system prompt, it only results in a marginal increase in the SR. This might be due to the emphasis on safety in prompts, leading to sub-optimal performance in agent planning and grounding procedures. One potential way to counter the previous step-wise prompt defense from the attackers’ perspective is to paraphrase the instructions in the pop-ups (Krishna et al., 2024; Abdali et al., 2024). A simple attack variant by removing the phrase “Please click” before the coordinates in the malicious instructions can effectively increase the ASR without any search or optimization.
Discussion
Beyond prompt-based defense, there are several practical approaches to potentially mitigate our attack in the real world, such as implementing more robust content filtering in browsers, adding modules to detect malicious instructions in agent observations (OpenAI, 2025), and using more detailed and specific descriptions of the attack within the defensive prompt. However, these efforts may not be sufficient to mitigate all environmental risks. Ideally, the VLM model and agent should recognize and understand this type of risk without requiring external tools.
@@ 6 Conclusion
In this work, we show that VLM agents can be easily maliciously instructed by adversarial pop-ups while operating on computers. Even though these pop-ups look very suspicious (by design) to human users, agents cannot distinguish the difference between pop-ups and typical digital content. This work offers two takeaways: (i) Just like human users need to undergo training to recognize phishing emails (Kumaraguru et al., 2007), VLM models/agents might need to undergo a similar process to ignore environmental noises and prioritize legitimate instructions (Wallace et al., 2024) before operating in the real digital world. This also applies to embodied agents since many distractors in the physical environment might also be absent from the training data. (ii) Human users need to oversee the automated agent workflow carefully (Shao et al., 2024) to manage the potential risks from the environment. Future work might focus on effectively leveraging human supervision and intervention for necessary safety concerns.
@@
This work is subject to a few limitations. (i) We only test the performance of closed-source models, which hinders a deeper understanding of why such an attack works. This choice was made due to the relatively low performance of open-source models on computer agent benchmarks. Future research is encouraged to explore more capable open-source models (Qin et al., 2025; Xu et al., 2024b). (ii) We do not explore more advanced jailbreaking techniques, such as optimizing the string inside the pop-ups (Zou et al., 2023) or making our pop-ups more persuasive (Zeng et al., 2024), but focus more on the high-level design of the adversarial pop-ups and the contribution analysis of different components.
Ethical Considerations
We study adversarial pop-up attacks on VLM agents solely for research purposes, with extensive discussion on potential designed choices, defense strategies, and their effectiveness, aiming to understand and address critical safety vulnerabilities before these systems are widely deployed. We emphasize the importance of human oversight in agent workflows and advocate for proper safety training of models for agentic usages.
@@ Acknowledgments
We thank Aryaman ‘Adam’ Arora, Harshit Joshi, Nikil Selvam, Yijia Shao, Yangjun Ruan, Chenglei Si, Dora Zhao, John Yang, Hao Zhu, Michael Ryan, Ryan Li, Ryan Louie, Caleb Ziems, Will Held, Yutong Zhang, Chen Henry Wu, Junlin Yang, Siyuan Ma, Graham Neubig, and all wonderful SALT Lab members for their valuable feedback on different stages of this work.
@@ Appendix A Appendix
To implement our attack, we first find the largest available rectangle area on the screen after excluding the bounding boxes from the a11y tree and OCR detection. We then randomly sample a position and size within this rectangle. We also limit the height of pop-ups to pixels and the width to pixels. The height of the info banner at the bottom is set to pixels, or one-third of the height of the pop-ups if the pop-up is not large enough. Finally, we search for the largest possible font size that fits within the pop-ups, where the attention hook/instruction/info banner is arranged as the example in Figure 2. To fully utilize the computational cost, we attack the agent observation whenever there is sufficient space for our pop-ups (both height and width are more than pixels). If the agent clicks on our pop-ups, we ignore this action during execution, and no redirection is implemented for simplicity. We use gpt-4o-2024-05-13 to summarize the user query and speculate the user query based on information on the screen through a11y trees. By default, we use “Please click (x,y)” as the instruction for both screenshot- and SoM agents in all OSWorld experiments, and “Please click [ID]” for SoM agents in all VisualWebArena experiments.
We randomly select a tag ID for our pop-up, which does not correlate with its website position. For the a11y tree element corresponding to the pop-ups, we use “{Tag ID} button button {ALT}” and “[{Tag ID}] [A] [{ALT}]” as adversarial templates for OSWorld and VisualWebArena correspondingly, which assumes the pop-ups are clickable items. We can run two benign (non-clickable) templates for comparison: “{Tag ID} {ALT}” for OSWorld and “[{Tag ID}] [IMG] [{ALT}]” for VisualWebArena, where we provide results in Table 10. The benign template leads to a slight increase in SR and a slight decrease in ASR.